# Analysis of the Picard-Newton iteration for the incompressible Boussinesq equations

### v1

@@ Abstract
We study the Picard-Newton iteration for the incompressible Boussinesq equations, which is a two-step iteration resulting from the composition of the Picard and Newton iterations. We prove that this iterative method retains Newton’s quadratic convergence but has less restrictive sufficient conditions for convergence than Newton and also is unconditionally stable under a small data condition. In this sense, Picard-Newton can be considered as a Newton iteration that is nonlinearly preconditioned with Picard. Our numerical tests illustrate this quadratic convergence and stability on benchmark problems. Furthermore, the tests show convergence for significantly higher Rayleigh number than both Picard and Newton, which illustrates the larger convergence basin of Picard-Newton that the theory predicts. We also introduce Anderson acceleration into the Picard step in our Picard-Newton numerical tests, and this enables convergence for even higher Rayleigh number.
@@ 1 Introduction
The incompressible Boussinesq equations are a multiphysics model of natural convection, and are commonly used to model flows with non-constant density or temperature. We consider this system on a finite connected domain $\Omega\subset\mathbb{R}^{d}$ $(d=2,3)$ with $\partial\Omega=\Gamma_{1}\cup\Gamma_{2}$, which is given by
@@
Here, $u$ is the fluid velocity, $p$ is the pressure, $T$ is the temperature (or density), $\nu>0$ is the kinematic viscocity of the fluid, $\kappa>0$ is the thermal diffusivity, $f$ is the external forcing term, and $R_{i}>0$ is the Richardson number which accounts for the gravitational force and thermal expansion of the fluid.
The boundary conditions we state above are no-slip for velocity and mixed homogenous Dirichlet and homogenous Neumann for temperature (the latter of corresponds to perfect insulation), but our results are extendable to most common boundary conditions. Important physical constants are the Reynolds number $Re$, Prandtl number $Pr$ which describes the ratio of kinematic viscocity and thermal diffusivity, and the Rayleigh number $Ra$:
@@
The Picard iteration and Newton iteration are common iterative methods for solving the Boussinesq equations. The Picard iteration is given by
@@
and is known to admit stable solutions for any problem data. Furthermore provided the data satisfies a smallness condition, the solution to Picard is unique and the iteration convergences linearly for any initial guess. Note that the system (3) decouples the equations because the heat transport equation is independent of velocity. While this all makes Picard very attractive with respect to efficiency at each iteration, the linear convergence is a potential drawback [15, 9, 24].
The Newton iteration for Boussinesq is given by
@@
with boundary conditions (4). This iteration will be shown in Section 2 to be well-posed and converge quadratically, but only for a very good initial guess and sufficiently small data (the sufficient conditions are more restrictive than for Picard). Newton is also more computationally expensive at each iteration, since it does not decouple and so one must solve fully coupled linear systems at each iteration. We note that it is common in fluid solvers to use Picard to find an initial guess for Newton by running Picard for several iterations.
The purpose of this paper is to consider the iteration defined by, at each step, first applying Picard and then applying Newton (called the Picard-Newton iteration). There are many other variations of utilizing both Picard and Newton together, which have improved efficiency and/or robustness of solvers for various problems [23, 14, 21, 16]. Our intent of combining the iterations in this way is to unite the speed of Newton with the stability and robustness of Picard. This method is simple to implement and we show it has an improved robustness compared to Newton but still converges quadratically. A similar improved robustness was shown both analytically and numerically in [18] for Picard-Newton applied to the Navier-Stokes equations.
We show herein analytically and numerically that the Picard-Newton iteration converges quadratically, with better convergence properties than Newton as well as improved stability. Due to our analytic results, the Picard-Newton iteration can be considered as the Newton iteration preconditioned with Picard, and thus fits a theme of recent and fundamentally important research called nonlinear preconditioning [5, 13, 12]. We also show this method can be further improved using Anderson-acceleration (AA) in the Picard Step, which we call AAPicard-Newton.
This paper is arranged as follows. We provide analysis of the well-posedness, stability, and convergence of Picard and Newton iterations in section 2. While much of this is known, the proofs and results will allow for a better comparison and analysis of Picard-Newton. In section 3 we give analysis on the stability and convergence of Picard-Newton. Following this in section 4, we provide numerical results for Picard-Newton applied to some benchmark tests. Within this section we also combine Picard-Newton with AA and give numerical results that reveal this gives further improvement.
@@ 2 Preliminaries
$|b(u,v,w)|$ $\leq C_{s}\|\nabla u\|\|\nabla v\|\|\nabla w\|$ (6)
$|\hat{b}(u,s,t)|$ $\leq C_{s}\|\nabla u\|\|\nabla s\|\|\nabla t\|$ (7)
@@
2.1 Boussinesq equations background
The weak form of the Boussinesq equations is: Given $f\in X^{\ast}$ and $g\in D^{\ast}$, find ($\text{find }(u,p,T)\in(X,Q,D)\text{ satisfying}$ $\forall(v,q,w)\in(X,Q,D)$,
@@ 2.1 Boussinesq equations background
Any solution to the Boussinesq equations (8) satisfies the a priori estimate
@@
The proof gives the bounds (9) and (10) for $T$ and $u$ respectively, and will be used throughout the paper for solutions to the Boussinesq equations. These results are similar to those found in [3], and depending on how the problem is formulated (e.g. Richardson number $\neq 1$ or Prandtl number $\neq 1$), the constants can differ but the results are equivalent. It is known that (8) admits solutions for any $Ra>0$ [3], and this can be proven in the same way as the steady Navier-Stokes existence proof using the Leray-Schauder theorem [11]. Uniqueness of (8), however, requires a smallness assumption on the data. Below is a sufficient condition on the problem data that produces uniqueness.
@@ 2.2 Picard iteration
The weak formulations for Picard for the Boussinesq system takes the form: Find $(u^{k+1},T^{k+1})\in V\times D$ satisfying $\forall(v,w)\in V\times D$,
@@
Any solution to the Picard iteration for the Boussinesq equations satisfies the a priori estimate: for any $k=1,2...$,
@@
Note that the sufficient conditions for convergence of Picard are the same as the sufficient conditions for the uniqueness of solutions to the Boussinesq equations.
@@
We subtract (8) from (13) and choose $v=e^{k+1}$ and $q=e_{T}^{k+1}$. Using skew-symmetry , vanishes two non-linear terms and leaves the equality
@@ 2.3 Newton iteration
The weak formulation for the Newton iteration for the Boussinesq system takes the form: Find $(u^{k+1},T^{k+1})\in V\times D$ satisfying $\forall(v,w)\in V\times D$,
@@
With this well-posedness result, the solution operator of (14) defined by $g_{N}:(V,D)\rightarrow(V,D)$, $g_{N}(u^{k},T^{k})=(u^{k+1},T^{k+1})$ is well-defined provided the data restrictions of Lemma 2.6 are satisfied. We note this is only a sufficient condition, and the Newton iteration for the Boussinesq equations (14) is believed to be well-posed on a much less restrictive set of parameters.
@@
3 Analysis of the Picard-Newton iteration
In this section, we consider the Picard-Newton iteration for the Boussinesq system, then analyze its stability and convergence properties. We show that the incorporation of a Picard step stabilizes the Newton step and reduces the (sufficient conditions on) small data assumptions and closeness assumption of the initial guess.
Using the solution operators for Picard and Newton defined in the previous section, the Picard-Newton iteration is defined as $g_{PN}:=g_{N}\circ g_{P}$ so that
@@ 3 Analysis of the Picard-Newton iteration
Thus, the Picard-Newton iteration consists of two steps. The first step is the Picard step given by $g_{P}(u^{k},T^{k})=(\tilde{u}^{k+1},\tilde{T}^{k+1})$ which is then followed by the Newton step given by $g_{N}(\tilde{u}^{k+1},\tilde{T}^{k+1})=(u^{k+1},T^{k+1})$.
@@
If $\alpha_{1}+\alpha_{2},~{}\nu^{-1}R_{i}C_{p}^{2}<1$ the Picard-Newton iteration (16)-(17) is stable with
@@
The small data conditions for stability of Picard-Newton are the same as the uniqueness condition for the Boussinesq equations.
@@
We begin by considering the stability bound resulting from the Picard step. This part of the proof follows analagously to Lemma 2.3 resulting in
@@
Assume $C_{p}^{2}Ri\nu^{-1}<1$. Then the Picard step (16) of the Picard-Newton iteration satisfies
@@
This proof follows analagously to Lemma 2.5.
@@
for any $k$. Then the Picard-Newton iteration will converge quadratically for any $u^{k}$ satisfying
@@
However there is another important connection to be made between Picard and the sufficient conditions. If we consider just Picard, it is required for its convergence that $\alpha_{1}+\alpha_{2}<1$. Then for Picard-Newton, $\alpha_{1}$ and $\alpha_{2}$ come from the error of the Picard step and appear in the sufficient conditions and assumptions of the method. This means that Picard-Newtons conditions for convergence are potentially being aided by Picard’s convergence properties. Furthermore, recall that the assumption for uniqueness of solutions to Boussinesq equations includes $\alpha_{1}+\alpha_{2}<1$. This further shows the assumption is reasonable and therefore Picard-Newton’s sufficient conditions for convergence of Picard-Newton are improved when compared to Newton.
@@ 4 Numerical Results
We give now numerical results for two test problems for Picard-Newton applied to the Boussinesq equations. We consider the Boussinesq system with $g=0$, $f=0$, $\nu=\kappa=10^{-1}$, and $Ri$ is varied which varies $Ra$ for the problems. The domain $\Omega$ and boundary conditions are different for each test and will therefore be provided in the following subsections.
@@
and we use 200 iterations as the maximum number of allowed iterations. We also use AA in the Picard step of Picard-Newton in some tests cases to illustrate possible further improvements.
@@ 4.1 Differentially Heated Cavity
We use a mesh with max element diameter $h=1/64$ and note that all results are comparable for other choices of h that we tested. Solutions to this system with $Ra=10000$ are shown in figure 1.
We study the convergence of Picard-Newton for varying $Ra$. Convergence results are shown in figure 2 for $Ra=50000$, $100000$, $150000$, $200000$, and $250000$. We observe that Picard-Newton converges for each $Ra$, and quadratic convergence is observed in the asymptotic range. As Ra increases, the number of iterations required for convergence increases as well. We observe the method remains stable even when it is (seemingly) not making progress in its residual.
Figure 1: Shown above is the computed Boussinesq solution of the differentially heated cavity problem for velocity streamlines (left) and temperature contours (right) for $Ra=10000$
For comparison, we also perform tests using Picard alone and Newton alone. In figure 3 (left) we see that Picard converges within 200 iterations for $Ra$ up to $10000$. For higher $Ra$, Picard does not converge, but it does remain stable. Newton is able to converge for slightly higher $Ra$ than Picard, as shown in figure 3 (left): it converges for up to $Ra=15000$, but above this $Ra$ we observe divergence. Moreover, as is typical for Newton, when it diverges, it blows up. The maximum Rayleigh number at which each method converges is given in table 2, and we observe that Picard-Newton is able to converge for $Ra$ a full order of magnitude higher than Picard and Newton.
Method Max $Ra$ for convergence
Picard-Newton $250000$
Picard $10000$
Newton $15000$
Table 2: Shown above is the maximum $Ra$ (to the nearest 1000) at which each method converges.

### v2

# Anderson-Picard based nonlinear preconditioning of the Newton iteration for non-isothermal flow simulations

@@ Abstract
We propose, analyze, and test a nonlinear preconditioning technique to improve the Newton iteration for non-isothermal flow simulations. We prove that by first applying an Anderson accelerated Picard step, Newton becomes unconditionally stable (under a uniqueness condition on the data) and its quadratic convergence is retained but has less restrictive sufficient conditions on the Rayleigh number and initial condition’s accuracy. Since the Anderson-Picard step decouples the equations in the system, this nonlinear preconditioning adds relatively little extra cost to the Newton iteration (which does not decouple the equations). Our numerical tests illustrate this quadratic convergence and stability on multiple benchmark problems. Furthermore, the tests show convergence for significantly higher Rayleigh number than both Picard and Newton, which illustrates the larger convergence basin of Anderson-Picard based nonlinear preconditioned Newton that the theory predicts.
@@ 1 Introduction
The non-isothermal flow simulations are a multiphysics model of natural convection, and are commonly used to model flows with non-constant density or temperature. We consider this system, often called the Boussinesq system, on a finite connected domain $\Omega\subset\mathbb{R}^{d}$ $(d=2,3)$ with $\partial\Omega=\Gamma_{1}\cup\Gamma_{2}$, and is given by
@@
Here, $u$ is the fluid velocity, $p$ is the pressure, $T$ is the temperature (or density), $\nu>0$ is the kinematic viscosity of the fluid, $\kappa>0$ is the thermal diffusivity, $f$ is the external forcing term, and $R_{i}>0$ is the Richardson number which accounts for the gravitational force and thermal expansion of the fluid. This model can be observed in nature such as in atmospheric models and katabatic winds, and is also observed in industry for natural ventilation, dense gas dispersion, insulation with double pane window, and solar collectors [6].
The boundary conditions we state above are no-slip for velocity and mixed homogenous Dirichlet and homogenous Neumann for temperature (the latter of which corresponds to perfect insulation), but our results are extendable to most common boundary conditions. Important physical constants are the Reynolds number $Re$, Prandtl number $Pr$ which describes the ratio of kinematic viscosity and thermal diffusivity, and the Rayleigh number $Ra$:
@@
The non-isothermal flow system (1)-(2) is known to be well-posed under a small data condition [3], however uniqueness is lost for high $Ra$.
The Picard iteration and Newton iteration are common iterative methods for solving the system (1)-(2). The Picard iteration for this system is given by
@@
and is known to admit stable solutions for any problem data (see Section 2). Furthermore, provided the data satisfies a smallness condition, the solution to Picard is unique and the iteration convergences linearly for any initial guess (see Section 2). Note that the system (3) decouples the equations because the heat transport equation is independent of velocity. While this all makes Picard very attractive with respect to efficiency at each iteration, the linear convergence is a potential drawback [15, 10, 20].
The Newton iteration for the system (1)-(2) is given by
@@
with boundary conditions (4). This iteration will be shown in Section 2 to be well-posed and converge quadratically, but only for a very good initial guess and sufficiently small data (the sufficient conditions are more restrictive than for Picard). Newton is also more computationally expensive at each iteration compared to Picard since it does not decouple, and so one must solve fully coupled linear systems at each iteration.
The purpose of this paper is to consider the iteration defined by, at each step, first applying Anderson accelerated Picard and then applying Newton (i.e. Anderson-Picard based nonlinearly preconditioned Newton). Our intent of combining the iterations in this way is to unite the speed of Newton with the stability and robustness of Picard. This method is simple to implement and we show it has an improved robustness compared to Newton (both with respect to stability and the size of the convergence basin), but still converges quadratically. This work thus fits a theme of recent and fundamentally important research on nonlinear preconditioning [5, 14, 13].
The general idea of applying Anderson-Picard as a preconditioner was first done in 2024 when it was applied to the Navier-Stokes Equations (NSE) [17]. The purpose of this paper is to extend this important yet easy to implement idea to non-isothermal flow. The difference between NSE and non-isothermal flow simulations is foremost that it is coupled with another equation which incorporates temperature (or density) into the model. This is further complicated by the extra nonlinear terms that arise in this coupled equation. This consequently gives rise to extra difficulties in the analysis and computation, especially for iterative methods like Newton where the equations remain fully coupled. In other words, it is not at all a straightforward extension. Moreover, the analysis provided in this paper will reveal that the convergence basin for Anderson-Picard-Newton is dependent upon only the initial conditions for velocity. Note that for non-preconditioned Newton it depends on both the initial conditions of velocity and temperature (or density). This is concurrent with the accuracy of Anderson-Picard-Newton depending more critically on velocity in a similar manner.
We show herein analytically and numerically that the (Anderson-)Picard based nonlinearly preconditioned Newton, shortened to (AA-)Picard-Newton, converges quadratically, with better convergence properties than Newton as well as improved stability. We restrict our analysis to Picard-Newton, i.e no Anderson, and note that the results are extendable to AA-Picard-Newton with some extra technical difficulties. We also provide numerical results for both AA-Picard-Newton and Picard-Newton that show a significant improvement compared to Newton alone.
This paper is arranged as follows. We provide analysis of the well-posedness, stability, and convergence of Picard and Newton iterations in section 2. While much of this is known, the proofs and results will allow for a better comparison and analysis of Picard-Newton. In section 3 we give analysis on the stability and convergence of Picard-Newton. Following this in section 4, we provide numerical results for both AA-Picard-Newton and Picard-Newton applied to some benchmark tests.
@@ 2 Preliminaries
$|b(u,v,w)|$ $\leq C_{s}\|\nabla u\|\|\nabla v\|\|\nabla w\|,$ (6)
$|\hat{b}(u,s,t)|$ $\leq C_{s}\|\nabla u\|\|\nabla s\|\|\nabla t\|.$ (7)
@@
2.1 Anderson-Acceleration
Anderson Acceleration (AA) is an extrapolation technique which utilizes previously computed iterates to construct the next solution. It is known to improve convergence for linearly convergent fixed point iterations like Picard [18, 8, 16, 19, 11, 8]. Therefore, it is a natural to precondition Newton with Anderson-Picard.
Algorithm 2.1.
AA for depth $m\geq 0$
Step 0: Choose $x_{0}\in Y$
Step 1: Find $w_{1}\in Y$ such that $w_{1}=g(x_{0})-x_{0}$. Set $x_{1}=x_{0}+w_{1}$.
Step k: For $k=2,3,...$. Set $m_{k}=\min\{k-1,m\}$.
  [a.] Find $w_{k}=g(x_{k-1})-x_{k-1}$.
  [b.] Solve the minimization problem for $\{\alpha_{j}^{k}\}_{k-m_{k}}^{k-1}$
  [c.] Set
where $w_{j}=g(x_{j-1})-x_{j-1}$ may be referred to as the update step or also as the nonlinear residual. Setting $m=0$ returns the original fixed point iteration.
2.2 Non-isothermal flow background
The weak form of the non-isothermal flow system (1)-(2) is: Given $f\in X^{\ast}$ and $g\in D^{\ast}$, find ($\text{find }(u,p,T)\in(X,Q,D)\text{ satisfying}$ $\forall(v,q,w)\in(X,Q,D)$,
@@ 2.2 Non-isothermal flow background
Any solution to the non-isothermal flow system (8) satisfies the a priori estimate
@@
The proof gives the bounds (9) and (10) for $T$ and $u$ respectively, and will be used throughout the paper for solutions to the non-isothermal flow system. These results are similar to those found in [3], and depending on how the problem is formulated (e.g. Richardson number $\neq 1$ or Prandtl number $\neq 1$), the constants can differ but the results are equivalent. It is known that (8) admits solutions for any $Ra>0$ [3], and this can be proven in the same way as the steady Navier-Stokes existence proof using the Leray-Schauder theorem [12]. Uniqueness of (8), however, requires a smallness assumption on the data. Below is a sufficient condition on the problem data that produces uniqueness.
@@ 2.3 Picard iteration
The weak formulations for Picard for the non-isothermal flow system takes the form: Find $(u^{k+1},T^{k+1})\in V\times D$ satisfying $\forall(v,w)\in V\times D$,
@@
Any solution to the Picard iteration for the non-isothermal flow system satisfies the a priori estimate: for any $k=1,2...$,
@@
Note that the sufficient conditions for convergence of Picard are the same as the sufficient conditions for the uniqueness of solutions to the non-isothermal flow system.
@@
We subtract (8) from (13) and choose $v=e^{k+1}$ and $q=e_{T}^{k+1}$. Using skew-symmetry , vanishes two nonlinear terms and leaves the equality
@@ 2.4 Newton iteration
The weak formulation for the Newton iteration for the non-isothermal flow system takes the form: Find $(u^{k+1},T^{k+1})\in V\times D$ satisfying $\forall(v,w)\in V\times D$,
@@
With this well-posedness result, the solution operator of (14) defined by $g_{N}:(V,D)\rightarrow(V,D)$, $g_{N}(u^{k},T^{k})=(u^{k+1},T^{k+1})$ is well-defined provided the data restrictions of Lemma 2.6 are satisfied. We note this is only a sufficient condition, and the Newton iteration for the non-isothermal flow system (14) is believed to be well-posed on a much less restrictive set of parameters.
@@
3 Nonlinear Preconditioner Analysis
In this section, we consider (AA-)Picard-Newton for the non-isothermal flow system, then analyze its stability and convergence properties. We show that preconditioning stabilizes Newton and reduces the sufficient conditions on small data assumptions and closeness assumption of the initial guess. We restrict our analysis to the case of no AA, however we note that with extra technicalities, similar proofs could be adapted.
Using the solution operators for Picard and Newton defined in the previous section, the Picard based nonlinear preconditioning of Newton, Picard-Newton, is defined as $g_{PN}:=g_{N}\circ g_{P}$ so that
@@ 3 Nonlinear Preconditioner Analysis
Thus, Picard-Newton consists of two steps. The first step is the Picard step given by $g_{P}(u^{k},T^{k})=(\tilde{u}^{k+1},\tilde{T}^{k+1})$ which is then followed by the Newton step given by $g_{N}(\tilde{u}^{k+1},\tilde{T}^{k+1})=(u^{k+1},T^{k+1})$.
@@
If $\alpha_{1}+\alpha_{2},~{}\nu^{-1}R_{i}C_{p}^{2}<1$, Picard-Newton (16)-(17) is stable with
@@
The small data conditions for stability of Picard-Newton are the same as the uniqueness condition for the non-isothermal flow system.
@@
We begin by considering the stability bound resulting from the Picard step. This part of the proof follows analogously to Lemma 2.3 resulting in
@@
Assume $C_{p}^{2}Ri\nu^{-1}<1$. Then the Picard step (16) of Picard-Newton satisfies
@@
This proof follows analogously to Lemma 2.5.
@@
for any $k$. Then Picard-Newton converge quadratically for any $u^{k}$ satisfying
@@
However there is another important connection to be made between Picard and the sufficient conditions. If we consider just Picard, it is required for its convergence that $\alpha_{1}+\alpha_{2}<1$. Then for Picard-Newton, $\alpha_{1}$ and $\alpha_{2}$ come from the error of the Picard step and appear in the sufficient conditions and assumptions of the method. This means that Picard-Newtons conditions for convergence are potentially being aided by Picard’s convergence properties. Furthermore, recall that the assumption for uniqueness of solutions to non-isothermal flow system includes $\alpha_{1}+\alpha_{2}<1$. This further shows the assumption is reasonable and therefore Picard-Newton’s sufficient conditions for convergence of Picard-Newton are improved when compared to Newton.
@@ 4 Numerical Results
We give now numerical results for two test problems for (AA-)Picard-Newton applied to the non-isothermal flow system. We consider the non-isothermal flow system with $g=0$, $f=0$, $\nu=\kappa=10^{-1}$, and $Ri$ is varied which varies $Ra$ for the problems. The domain $\Omega$ and boundary conditions are different for each test and will therefore be provided in the following subsections.
@@
and we use 200 iterations as the maximum number of allowed iterations.
@@ 4.1 Differentially Heated Cavity
We use a mesh with max element diameter $h=1/64$ and note that all results are comparable for other choices of h that we tested. Solutions to this system with $Ra=10000$ are shown in figure 1. We begin the study of the convergence of (AA-)Picard-Newton for varying $Ra$ by considering Picard-Newton and then AA-Picard-Newton.
4.1.1 Picard-Newton
Convergence results of Picard-Newton are shown in figure 2 for $Ra=50000$, $100000$, $150000$, $200000$, and $250000$. We observe that Picard-Newton converges for each $Ra$, and quadratic convergence is observed in the asymptotic range. As Ra increases, the number of iterations required for convergence increases as well. We observe the method remains stable even when it is (seemingly) not making progress in its residual.
Figure 1: Shown above is the computed solution of the differentially heated cavity problem for velocity streamlines (left) and temperature contours (right) for $Ra=10000$
For comparison, we also perform tests using Picard alone and Newton alone. In figure 3 (left) we see that Picard converges within 200 iterations for $Ra$ up to $10000$. For higher $Ra$, Picard does not converge, but it does remain stable. Newton is able to converge for slightly higher $Ra$ than Picard, as shown in figure 3 (left): it converges for up to $Ra=15000$, but above this $Ra$ we observe divergence. Moreover, as is typical for Newton, when it diverges, it blows up. We observe then that that Picard-Newton is able to converge for $Ra$ a full order of magnitude higher than Picard and Newton.